# RBI-ThPhys-2025-26

### v2

@@ Abstract
We measure three-dimensional bispectra of halo intrinsic alignments (IA) and dark matter overdensities in real space from N-body simulations for halos of mass $10^{12}-10^{12.5}M_{\odot}/h$. We show that their multipoles with respect to the line of sight can be accurately described by a tree-level perturbation theory model on large scales ($k\lesssim 0.11\,h$/Mpc) at $z=0$. For these scales and in a simulation volume of 1 Gpc/$h$, we detect the bispectrum monopole $B_{\delta\delta E}^{00}$ at $\sim 30\sigma$ and the two quadrupoles $B_{\delta\delta E}^{11}$ and $B_{\delta\delta E}^{20}$ at $\sim 25\sigma$ and $\sim 15\sigma$, respectively. We also report similar detection significances for the lowest order multipoles of $B_{\delta EE}$ and $B_{EEE}$, although these are largely driven by stochastic contributions. We show that the first and second order EFT parameters are consistent with those obtained from fitting the IA power spectrum analysis at next-to-leading order, without requiring any priors to break degeneracies for the quadratic bias parameters. Moreover, the inclusion of higher multipole moments of $B_{\delta\delta E}$ greatly reduces the errors on second order bias parameters, by factors of 5 or more. The IA bispectrum thus provides an effective means of determining higher order shape bias parameters, thereby characterizing the scale dependence of the IA signal. We also detect parity-odd bispectra such as $B_{\delta\delta B}$ and $B_{\delta EB}$ at $\sim 10\sigma$ significance or more for k<0.15hk<0.15\,h/Mpc and they are consistent with the parity-even sector. Furthermore, we check that the Gaussian covariance approximation works reasonably well on the scales we consider here. These results lay the groundwork for using the bispectrum of IA in cosmological analyses.
@@ 1 Introduction
In stark contrast to the (primary) anisotropies of the cosmic microwave background, the distribution of LSS is highly non-Gaussian due to nonlinear gravitational evolution of the initial density perturbations. Hence, the statistics of LSS are not captured by only the two-point correlation function and linear theory does not suffice except at very large separations. This has triggered a surge in the study of statistics of LSS beyond the power spectrum, the simplest of which is the bispectrum, i.e. the Fourier transform of the three-point correlation function (6). In the case of correlations between galaxy positions, the addition of such non-Gaussian statistics has been shown to yield a gain in constraining power on the cosmological model parameters (7; 8; 9; 10; 11).
@@
The modelling of three-point statistics of weak gravitational lensing has been considered in numerous other works (see e.g. (15; 16; 17; 18) for recent applications). However, the contribution of IA to the observed three-point signal has received less attention (though see (19)) than the two-point IA signal, which has been extensively studied in e.g. (20; 21; 22). It is of note that this contribution may contain valuable information on interesting physical phenomena such as anisotropic local primordial non-Gaussianity (23; 24), in analogy to the galaxy bispectrum.
The objective of this paper and its earlier companion (25) (hereafter Paper I) is to assess the gain in information on shape statistics by including the IA bispectrum. On the modelling side, we restrict ourselves to large scales where the bispectrum signal can be reliably computed using perturbation theory. In addition, we will assume that three-dimensional positions of galaxies are available through spectroscopic redshift measurements. In that case, the bispectrum of intrinsic alignment can be summarized in terms of multipole moments with respect to the line of sight, for which Paper I developed separable estimators. Notably, such estimators are applicable regardless of the scales being modelled (as long as a flat-sky treatment is appropriate). In Paper I, we forecasted an SNR of $\approx 30$ for $B_{DDE}$ (where $D$ denotes the galaxy density) in a DESI-like LRG sample, where shapes are measured using LSST. Direct measurements of intrinsic alignments for samples with spectroscopic redshifts can aid in isolating this effect from cosmic shear (26; 27).
@@ 2 Tree-level Shape Bispectrum
The tree-level tensor bispectrum in the Effective Field Theory for Intrinsic Alignments (EFT of IA) was analyzed in Paper I. For this reason, we refer the reader to this paper for all the relevant details and restrict ourselves to a summary of its results (see also (31; 32; 28) for more details).
@@
As in Paper I, we did not consider higher derivative contributions to $g_{ij}$ (or the matter density field). Assuming that the non-locality scale of halos is similar (or smaller) than the nonlinear scale, such contributions would be of higher order than the ones we include in the tree-level bispectrum. We also did not consider correlations with scalar biased tracers in this paper.
@@ 2.1 Tree-level Bispectrum
$S_{ij}(\mathbf{x})=\frac{1}{3}\delta_{ij}\delta(\mathbf{x})+g_{ij}(\mathbf{x}),$ (4)
@@
We then obtain the mixed scalar ($\delta$) and tensor (g) bispectra by taking trace and trace-free parts of the full tensorial bispectrum:
@@
The three second order IA parameters are simply a redefinition of the more commonly used ‘TATT’ parameters $b_{KK},b_{t},b_{\delta K}$ (35; 28). It turns out that second order IA contributions are generally smaller than the linear alignment term proportional to (powers of) $b_{1}^{\text{g}}$ [^1] Indeed, (36) find that a simpler alignment model is sufficient to describe three-point alignments in the BOSS LOWZ sample.. However, as we will see, the halo sample considered here cannot be modelled without second order IA contributions. The leading stochastic shape bispectrum $\langle\epsilon_{ij}\epsilon_{kl}\epsilon_{rs}\rangle\propto\alpha_{\text{ggg}}^{111}$ turns out to vanish after projecting the tensor bispectrum onto the sky (see App. A of Paper I), which is the part we discuss next.
@@ 2.2 Flat Sky Projection
Not all five degrees of freedom of the shape field can be discerned by an observer at a fixed location. Instead, only two degrees of freedom in $g_{ij}$ are observable. They constitute a spin-2 field on the sky which is commonly decomposed into E- and B-modes in Fourier space. Throughout this work we will use the flat-sky approximation to construct projected observables. To recap the findings of (31; 32), the projection on the sky is performed by applying the trace-free operator
@@
The E- and B-modes can then be obtained by applying this projector to $g_{ij}(\mathbf{x})$, since it commutes with the flat-sky projection operator $\mathcal{P}_{ijkl}$ (31; 28):
@@
$\mathbf{M}_{ij}^{\delta}(\hat{\mathbf{k}})=\delta_{ij}$ (19)
@@
the spectra with two B-modes (that is, $B_{\delta BB}$ and $B_{EBB}$) necessarily only depend on the stochastic parameter $\alpha_{\text{gg}}^{1\delta}$; they do not depend on the $b_{2,a}^{\text{g}}$. The bispectrum consisting of three B-modes, $B_{BBB}$ vanishes identically in our model since the contribution from $\alpha_{\text{gg}}^{1\delta}$ also gets projected out. Furthermore, the deterministic piece of spectra containing any B-modes is typically strongly suppressed, and as such we do not expect these spectra to meaningfully constrain the $b_{2,a}^{\text{g}}$ bias parameters. For that reason, we will not consider the parity-odd bispectra $B_{\delta\delta B}$ and $B_{\delta EB}$ in any baseline analysis; instead we will only focus on fitting parity-even bispectra and subsequently check that the obtained results are broadly consistent with the parity-odd sector.
@@ 2.3 Bispectrum Multipoles
The bispectrum of tensor fields admits a multipole decomposition as detailed in Paper I. There it was argued that a separable local plane-parallel Yamamoto estimator requires the use of associated Legendre polynomials to define the multipole moments (38; 39). In that work, it was also shown that differences between the use of ordinary and associated Legendre polynomials are small and that associated Legendre polynomials appear better suited for $B_{\delta\delta E}$ (at least in real space, which is the case we restrict to here). For that reason, we will only consider associated Legendre multipoles (ALMs) here. We thus also omit the superscript ‘A’ on the multipoles that was present in Paper I to distinguish between ordinary and associated Legendre multipoles, and simply refer to the ALMs as multipoles. Specifically, defining the even and odd projectors
@@ 3 Simulations and Data Analysis
We will use a suite of 20 high-resolution dark-matter only N-body simulations from the DarkQuest Project (41). Each realization has $2048^{3}$ N-body particles with mass $M_{p}=1.02\times 10^{10}M_{\odot}/h$ and box side length $L=1\text{ Gpc}/h$. For more specifics, see e.g. (20). In this setup, we will take the halo sample in mass range $10^{12}-10^{12.5}M_{\odot}/h$ at redshift $z=0$, which we also analyzed in (28). This is especially convenient since it allows us to check whether bias parameters that are shared between the power spectrum and bispectrum of IA match. We defer performing a joint analysis to potential future work. The sample of halos in a single realization has different properties from the Luminous Red Galaxy (LRG) sample considered in Paper I. Notably, the number density is roughly $\bar{n}=3\times 10^{-3}(h/\text{Mpc})^{3}$ and $A_{\text{IA}}\approx 8$ compared to $\bar{n}=5\times 10^{-4}(h/\text{Mpc})^{3}$ and the alignment amplitude (defined in Eq. (30)) $A_{\text{IA}}\approx 4$ in Paper I, meaning that the signal-to-noise ratio (SNR) for a given total volume is much higher. This allows us to perform a more stringent test of the model under consideration.
@@ 3.1 Estimator
when $XYZ$ is parity-even. To this end, for the parity-even sector we will consider only ordered triangle bins which satisfy $k_{1}\geq k_{2}\geq k_{3}$. When enumerating these bins, they are ordered ‘lexicographically’, i.e. the first few bins are $(1,1,1),(2,1,1),(2,2,1),(2,2,2),(3,1,1),\dots$ et cetera. Note that this setup includes open triangles such as $(2,1,1)$ for which the bin centers do not form a closed triangle; we discuss this in more detail below. We use a radial bin width of $\Delta k=0.4/30\,h$/Mpc which approximately equals $2k_{f}$, where $k_{f}$ is the fundamental mode. For example, for triangles with side lengths up to the maximum wavenumber $k_{\text{max}}=0.2\,h/\text{Mpc}$ there are $15$ radial wavenumber bins, resulting in a total of $428$ ordered triangle bins. For the parity-odd bispectra, which have an odd number of B-modes, i.e. $B_{\delta\delta B},B_{\delta EB},B_{EEB}$ and $B_{BBB}$ we will consider all triangles with the sole restriction $k_{1}>k_{2}$; this results in $917$ triangles in the parity-odd sector for $k_{\text{max}}=0.2\,h/\text{Mpc}$. Note that this restriction must necessarily be applied for all bispectra except for $B_{\delta EB}$, since they are symmetric in their first two arguments. In the case of $B_{\delta EB}$, we add to these triangles also the ones with $k_{1}=k_{2}=k_{3}$, of which there are 15. Under these assumptions, the Gaussian covariance of all multipoles (both even and odd) is diagonal in the triangle bins [^2] In the even case, this was already argued in Paper I. For the odd case, it can be shown from Eq. (E9) from Paper I by observing that $P_{XB}(\mathbf{k})=0$ whenever $X\neq B$. The parity-odd covariance would not be diagonal if we also included the $k_{1}>k_{2}$ part of the data vector for $B_{\delta EB}$. and also has no cross-correlation between even and odd multipoles.
We will not include parity-odd bispectra in the likelihood at any point, since they are detected at small significance compared to the parity-even ones in a single simulation volume. However, we emphasize that these spectra are clearly nonzero and as we will show, their amplitudes are fully consistent with what is inferred from fitting the parity-even sector.
@@ 3.2 Covariance
and $P_{\text{NL}}$ is computed using the CCL wrapper for HaloFit (42; 43). In the expression for $b_{1}^{\text{g}}$, $D(z)$ is the linear growth factor normalized to $1$ today and $C_{1}\rho_{\text{cr}}=0.0134$ (14; 44). For the Gaussian covariance, we use the fiducial set of values
@@
Figure 1: Ratio of diagonal sample covariance to the theoretical Gaussian prediction for several different multipoles. The grey shade indicates the region where $k_{\text{max}}<0.05\,h/\text{Mpc}$; deviations in this region are likely due to discretization effects and do not indicate model failure.
@@
The comparison of Eq. (32) with the theoretical Gaussian covariance estimate is shown in Figure 1 for five different multipoles involving the $\delta$ and $E$ fields. For several increasing values of $k_{\text{max}}$, we consider the ratio of the sample and Gaussian covariances for all triangles with all sides below $k_{\text{max}}$ and plot the mean and sample variance of these ratios to reduce scatter. It is clear that on large scales (k<0.10hk<0.10\,h/Mpc), the Gaussian approximation works well, while for larger $k_{\text{max}}$ the Gaussian covariance (using the nonlinear matter power spectrum) tends to somewhat overpredict the error. This is likely due to our choice of using the nonlinear power spectrum rather than the linear one. Moreover, we observe at most $\sim 20\%$ offsets between the theoretical, Gaussian covariance and Eq. (32), which amounts to $\sim 10\%$ on the corresponding errors. We consider this sufficient for the purpose of this work.

### v3

@@ Abstract
[…]
@@ 1 Introduction
In stark contrast to the (primary) anisotropies of the cosmic microwave background, the distribution of LSS is highly non-Gaussian due to nonlinear gravitational evolution of the initial density perturbations. Hence, the statistics of LSS are not captured by only the two-point correlation function and linear theory does not suffice except at very large separations. This has triggered a surge in the study of statistics of LSS beyond the power spectrum, the simplest of which is the bispectrum, i.e. the Fourier transform of the three-point correlation function (6). In the case of correlations between galaxy positions, the addition of the bispectrum has been shown to yield a gain in constraining power on the cosmological model parameters. (7; 8; 9; 10; 11).
@@
The modeling of higher-order statistics of weak gravitational lensing has been considered in numerous other works (see e.g. (15; 16; 17; 18) for recent applications of three-point statistics, or (19; 20) for more general higher-order statistics). However, the contribution of IA to the observed three-point signal has received less attention (though see (21; 22)) than the two-point IA signal, which has been extensively studied in e.g. (23; 24; 25). It is of note that this contribution may contain valuable information on interesting physical phenomena such as anisotropic local primordial non-Gaussianity (26; 27), in analogy to the galaxy bispectrum.
The objective of this paper and its earlier companion (28) (hereafter Paper I) is to assess the gain in information on shape statistics by including the IA bispectrum. On the modelling side, we restrict ourselves to large scales where the bispectrum signal can be reliably computed using perturbation theory. The general perturbative framework developed in (29; 30; 31) (see also (32)) offers a systematic improvement over simpler models of IA, such as the ‘NLA’ bispectrum model used in (21). As we will see in this work, the higher-order corrections to the IA bispectrum are clearly present for the halo sample we consider and cannot be neglected. In addition, we will assume that three-dimensional positions of galaxies are available through spectroscopic redshift measurements. In that case, the bispectrum of intrinsic alignment can be summarized in terms of multipole moments with respect to the line of sight, for which Paper I developed separable estimators. These estimators capture the full scale dependence and angular structure of the bispectrum and do not restrict to specific configurations such as equilateral triangles or triangles in the sky plane. Notably, such estimators are applicable regardless of the scales being modelled (as long as a flat-sky treatment is appropriate). In Paper I, we forecasted an SNR of $\approx 30$ for $B_{DDE}$ (where $D$ denotes the galaxy density) in a DESI-like LRG sample, where shapes are measured using LSST. Direct measurements of intrinsic alignments for samples with spectroscopic redshifts can aid in isolating this effect from cosmic shear (33; 34).
@@ 2 Tree-level Shape Bispectrum
The tree-level tensor bispectrum in the Effective Field Theory for Intrinsic Alignments (EFT of IA) was analyzed in Paper I. For this reason, we refer the reader to this paper for all the relevant details and restrict ourselves to a summary of its results (see also (29; 30; 31) for more details).
@@
As in Paper I, we did not consider higher derivative contributions to $g_{ij}$ (or the matter density field). Assuming that the non-locality scale of halos is similar to (or smaller than) the nonlinear scale, such contributions would be of higher order than the ones we include in the tree-level bispectrum. We also did not consider correlations with scalar biased tracers in this paper.
@@ 2.1 Tree-level Bispectrum
$S_{ij}(\mathbf{x})=\frac{1}{3}\delta_{ij}^{\mathrm{K}}\delta(\mathbf{x})+g_{ij}(\mathbf{x}),$ (4)
@@
Here $\delta(\mathbf{x})$ is the matter overdensity and $\delta_{ij}^{\mathrm{K}}$ is the Kronecker delta. We then obtain the mixed scalar ($\delta$) and tensor (g) bispectra by taking trace and trace-free parts of the full tensorial bispectrum:
@@
The three second order IA parameters are simply a redefinition of the more commonly used ‘TATT’ parameters $b_{KK},b_{t},b_{\delta K}$ (39; 31). It turns out that second order IA contributions are generally smaller than the linear alignment term proportional to (powers of) the linear shape bias  [^1] Indeed, (22) find that a simpler alignment model is sufficient to describe three-point alignments in the BOSS LOWZ sample.. However, as we will see, the halo sample considered here cannot be modelled without second order IA contributions. The leading stochastic shape bispectrum $\langle\epsilon_{ij}\epsilon_{kl}\epsilon_{rs}\rangle\propto\alpha_{\text{ggg}}^{111}$ turns out to vanish after projecting the tensor bispectrum onto the sky (see App. A of Paper I), which is the part we discuss next.
@@ 2.2 Flat Sky Projection
Not all five degrees of freedom of the shape field can be discerned by an observer at a fixed location. Instead, only two degrees of freedom in $g_{ij}$ are observable. They constitute a spin-2 field on the sky which is commonly decomposed into E- and B-modes in Fourier space. Throughout this work we will use the flat-sky approximation to construct projected observables. For an in-depth study of the validity of the flat-sky approximation for various tracers, see e.g. (40). Note that current surveys may require wide-angle effects to be taken into account (41) - we leave this for future work. To recap the findings of (29; 30), the projection on the sky is performed by applying the trace-free operator
@@
The E- and B-modes can then be obtained by applying this projector to $g_{ij}(\mathbf{x})$, since it commutes with the flat-sky projection operator $\mathcal{P}_{ijkl}$ (29; 31):
@@
$\mathbf{M}_{ij}^{\delta}(\hat{\mathbf{k}})=\delta_{ij}^{\mathrm{K}}$ (19)
@@
the spectra with two B-modes (that is, $B_{\delta BB}$ and $B_{EBB}$) necessarily only depend on the stochastic parameter $\alpha_{\text{gg}}^{1\delta}$; they do not depend on the $b_{2,a}^{\text{g}}$. The bispectrum consisting of three B-modes, $B_{BBB}$ vanishes identically in our model since the contribution from $\alpha_{\text{gg}}^{1\delta}$ also gets projected out, i.e. it vanishes when applying the B-mode projector (28). Furthermore, the deterministic piece of spectra containing any B-modes is typically strongly suppressed, and as such we do not expect these spectra to meaningfully constrain the $b_{2,a}^{\text{g}}$ bias parameters. For that reason, we will not consider the parity-odd bispectra $B_{\delta\delta B}$ and $B_{\delta EB}$ in any baseline analysis; instead we will only focus on fitting parity-even bispectra and subsequently check that the obtained results are broadly consistent with the parity-odd sector.
@@ 2.3 Bispectrum Multipoles
The bispectrum of tensor fields admits a multipole decomposition as detailed in Paper I. There it was argued that a separable local plane-parallel Yamamoto estimator requires the use of associated Legendre polynomials to define the multipole moments (43; 44). In that work, it was also shown that differences between the use of ordinary and associated Legendre polynomials are small and that associated Legendre polynomials appear better suited for $B_{\delta\delta E}$ (at least in real space, which is the case we restrict to here). For that reason, we will only consider associated Legendre multipoles (ALMs) here. We thus also omit the superscript ‘A’ on the multipoles that was present in Paper I to distinguish between ordinary and associated Legendre multipoles, and simply refer to the ALMs as multipoles. Specifically, defining the even and odd projectors
@@ 3 Simulations and Data Analysis
We will use a suite of 20 high-resolution dark-matter only N-body simulations from the DarkQuest Project (46). Each realization has $2048^{3}$ N-body particles with mass $M_{p}=1.02\times 10^{10}M_{\odot}/h$ and box side length $L=1\text{ Gpc}/h$. For more specifics, see e.g. (23). In this setup, we will take the halo sample in mass range $10^{12}-10^{12.5}M_{\odot}/h$ at redshift $z=0$, which we also analyzed in (31). This is especially convenient since it allows us to check whether bias parameters that are shared between the power spectrum and bispectrum of IA match. We defer performing a joint analysis to potential future work. The sample of halos in a single realization has different properties from the Luminous Red Galaxy (LRG) sample considered in Paper I. Notably, the number density is roughly $\bar{n}=3\times 10^{-3}(h/\text{Mpc})^{3}$ and $A_{\text{IA}}\approx 8$ compared to $\bar{n}=5\times 10^{-4}(h/\text{Mpc})^{3}$ and the alignment amplitude (defined in Eq. (30)) $A_{\text{IA}}\approx 4$ in Paper I, meaning that the signal-to-noise ratio (SNR) for a given total volume is much higher. This allows us to perform a more stringent test of the model under consideration. Higher alignment amplitudes of $A_{\text{IA}}\sim 15$ arise for galaxy clusters, although such samples would have much lower number density than LRGs (47).
@@ 3.1 Estimator
when $XYZ$ is parity-even. This renders the covariance matrix diagonal in triangle space in the Gaussian approximation (28). To this end, for the parity-even sector we will consider only ordered triangle bins which satisfy $k_{1}\geq k_{2}\geq k_{3}$. When enumerating these bins, they are ordered ‘lexicographically’, i.e. the first few bins are $(1,1,1),(2,1,1),(2,2,1),(2,2,2),(3,1,1),\dots$ et cetera. Note that this setup includes open triangles such as $(2,1,1)$ for which the bin centers do not form a closed triangle; we discuss this in more detail below. We use a radial bin width of $\Delta k=0.4/30\,h$/Mpc which approximately equals $2k_{f}$, where $k_{f}=2\pi/L$ is the fundamental mode. The bin width is chosen to strike a balance between having too few modes per bin (and hence very large discreteness effects) and loss of information due to over-compressing the data vector. For example, for triangles with side lengths up to the maximum wavenumber $k_{\text{max}}=0.2\,h/\text{Mpc}$ there are $15$ radial wavenumber bins, resulting in a total of $428$ ordered triangle bins. For the parity-odd bispectra, which have an odd number of B-modes, i.e. $B_{\delta\delta B},B_{\delta EB},B_{EEB}$ and $B_{BBB}$ we will consider all triangles with the sole restriction $k_{1}>k_{2}$; this results in $917$ triangles in the parity-odd sector for $k_{\text{max}}=0.2\,h/\text{Mpc}$. Note that this restriction must necessarily be applied for all bispectra except for $B_{\delta EB}$, since they are symmetric in their first two arguments. In the case of $B_{\delta EB}$, we add to these triangles also the ones with $k_{1}=k_{2}=k_{3}$, of which there are 15. Under these assumptions, the Gaussian covariance of all multipoles (both even and odd) is diagonal in the triangle bins [^2] In the even case, this was already argued in Paper I. For the odd case, it can be shown from Eq. (E9) from Paper I by observing that $P_{XB}(\mathbf{k})=0$ whenever $X\neq B$. The parity-odd covariance would not be diagonal if we also included the $k_{1}>k_{2}$ part of the data vector for $B_{\delta EB}$. and also has no cross-correlation between even and odd multipoles.
We will not include parity-odd bispectra in the likelihood at any point, since they are detected at small significance compared to the parity-even ones in a single simulation volume. However, we emphasize that these spectra are clearly nonzero and as we will show, their amplitudes are fully consistent with what is inferred from fitting the parity-even bispectra.
@@ 3.2 Covariance
and $P_{\text{NL}}$ is computed using the CCL wrapper for HaloFit (48; 49). In the expression for $b_{1}^{\text{g}}$, $D(z)$ is the linear growth factor normalized to $1$ today and $C_{1}\rho_{\text{cr}}=0.0134$ (14; 50). For the Gaussian covariance, we use the fiducial set of values
@@
Figure 1: Ratio of diagonal sample covariance to the theoretical Gaussian prediction for several different multipoles. The grey shade indicates the region where $k_{\text{max}}<0.05\,h/\text{Mpc}$ (this region was chosen somewhat artifically); deviations in this region are likely due to discretization effects and do not indicate model failure.
@@
The comparison of Eq. (32) with the theoretical Gaussian covariance estimate is shown in Figure 1 for five different multipoles involving the $\delta$ and $E$ fields. For several increasing values of $k_{\text{max}}$, we consider the ratio of the sample and Gaussian covariances for all triangles with all sides below $k_{\text{max}}$ and plot the mean and sample variance of these ratios to reduce scatter. On large scales (k<0.10hk<0.10\,h/Mpc), the Gaussian approximation works well (aside from very large scales k<0.05hk<0.05\,h/Mpc, where discreteness effects occur in the covariance matrix), while for larger $k_{\text{max}}$ the Gaussian covariance (using the nonlinear matter power spectrum) tends to somewhat overpredict the error. Moreover, we observe at most $\sim 20\%$ offsets between the theoretical, Gaussian covariance and Eq. (32), which amounts to $\sim 10\%$ on the corresponding errors. We consider this sufficient for the purpose of this work. We do not use any large-scale cuts (i.e. $k_{\text{min}}$) in the remainder of the paper, since large scales do not impact constraints significantly.